Fill in the boxes with the correct symbol out of >,
-5/7 [ ] 2/3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 7 and 3 is 21
∴(-5/7) = [(-5 × 3)/ (7 × 3)] = (-15/21)
And (2/3) = [(2 × 7)/ (3 × 7)]
= (14/21)
Now,
-15 < 14
So, (-15/21) < (14/21)
Hence, -5/7 [<] 2/3

Thanks